OpenStudy (anonymous):

a total of $2250 for a house painting job is to be divided between two painters in the ratio of 2 to 3 .how much does each painter receive ? thanks a lot ♥

OpenStudy (anonymous):

are you in alg 2? just a guess, lol it's all about setting up the equation - the rest is just following procedures 2x = 3y x+y=2250 blah x = 1350 y = 900

OpenStudy (anonymous):

1st year college .. college algebra

OpenStudy (anonymous):

as above, use 2x = 3y, and x+y=2250.... make x=2250-y and substitute this into 2x=3y to solve for y. once you have this number, then you sub it back in again and solve for x :)

OpenStudy (anonymous):

the first painter will get $900 and the second painter will get $1350 total for the ratio is 2+3=5 so $2250\5=$450 each person the ratio for first painter is 2 then $450x2=$900 the ratio for second painter is 3 then $450x3=$1350

OpenStudy (anonymous):

thanks a lot